OpenStudy (anonymous):

Select all that apply. Solve for x, 0 <_ x <_ 2(pi) 2sinx - sqrt(3) = 0 ___ (pi)/3 ___(2pi)/3 ___(4pi)/3 ___(5pi)/3 ** <_ means less than or equal to. I kind of need step by step because I really don't know what I'm doing here.. I also have more if someone is willing to continue to help me :) Thanks so much!

OpenStudy (anonymous):

start with \[2\sin(x)=\sqrt{3}\] then go to \[\sin(x)=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

now you have to look for angles that lie within the range given to you that satisfy this equation

OpenStudy (anonymous):

then find the angle (number) whose sine is \(\frac{\sqrt{3}}{2}\) which you can either memorize, or look at the unit circle on the last page of the attached cheat sheet.

OpenStudy (anonymous):

the the first coordinate on the unit circle is cosine, the second coordinate is sine find the corresponding angle. you should see that one of them is \(x=\frac{\pi}{3}\)

OpenStudy (anonymous):

Oh okay! That was a lot easier than I had imagined.. I see now, thank you :) I really appreciate it.

OpenStudy (anonymous):

there is another one also, which you should be able to see from the cheat sheet

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

The other is 2pi/3 :)

OpenStudy (anonymous):

@satellite73 I have a quick question.. If I have to sqrt something, like 1/4. I would get (+/-1)/(+/-2) Right? Well should I take 1/2 & -1/2 and find the corresponding angle? If not, how do I know if I take the negative or positive?